if a 0.12M solution of an acid is 4.2% ionized, calculate the pH of the solution and the dissociation constant of the acid
1
Expert's answer
2020-10-12T13:31:32-0400

Solution.

"Ka = \frac{\alpha^2 \times C}{1-\alpha}"

"Ka = 2.21 \times 10^{-4}"

"pH = -lg(\alpha \times C) = 2.30"

Answer:

"pH = 2.30"

"Ka = 2.21 \times 10^{-4}"